[figure: 9.7 cm]

STRUCTURE, TOLERANT FACTOR AND PROPERTIES OF (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} MAGNETORESISTIVE MATERIALS

Pashchenko V.P., Shemyakov A.A., Prokopenko V.P., Bybik O.B., D’yakonov V.P., ^{∗}Szymczak H.

Donetsk Physico-Technical Institute of Ukrainian NAS, 340114, Donetsk, Ukraine ∗Institute of Physics PAN, 02-668 Warszawa, Poland

The structure and properties of (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} (0 ≤ x ≤ 0.3) single-phase ceramic samples were investigated by X-rays, NMR on ^{55}Mn, magnetoelectical methods. Tolerant factor (t) calculated for ideal and real radiuses of ideal and defect lattices was monotonous decreased at x increasing. Wide NMR spectra^{55}Mn testified about different valent states of ions Mn^{3+}-Mn^{4+}. There is a high-frequency electron exchange between them. The amplitude of NMR spectrum practically was not depend on composition, and spectrum width, ΔF, was monotonous decreased at x increasing. At x increasing the temperature of magnetoresistance peak, ΔR/R_{0}, measured under 5 kOe, was not changed practically (T_{p}=265 K). Magnetoresistance at that was grown 2 times. The correlation between x, t_{r}, ΔF and ΔR/R_{0} was most clearly displayed at the account of real ion radiuses, moreover, with different coordination number (12, 9, 8, 6) of defect lattice contained simultaneously cationic and anionic vacancies.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr. Pashchenko V.P.

Address for correspondence : Donetsk Physico-Technical Institute, R. Luxembourg str., 72, 340114, Donetsk, Ukraine

Email address : pashchen@host.dipt.donetsk.ua